OpenStudy (anonymous):

The downtown bus leaves the station leaves the bus station every 15 min. The upstown bus leaves the station every 35 min. Both buses start the day leaving the station at 7:00 A.M. When is the next time both buses will leave the station at the same time?

OpenStudy (anonymous):

15 = 3 *5 35 = 5 * 7 lcm = 3*5*7 = 105 minutes 7 am +105 minutes = ....

OpenStudy (anonymous):

Uhhhhh

OpenStudy (anonymous):

That makes 112 min. but what time.... O.o

OpenStudy (anonymous):

convert minutes to hour 1 hour = 60 minutes

OpenStudy (anonymous):

Ok so 8:12?

OpenStudy (anonymous):

how you got 112?

OpenStudy (anonymous):

Idk.... Lol

OpenStudy (anonymous):

7+105....

OpenStudy (anonymous):

not like that 105 minutes = (105/60) hour = 1.75 hour = 2 hour 15 minutes 7 am + 2 hour 15 minutes = ...

OpenStudy (anonymous):

142

OpenStudy (anonymous):

??

OpenStudy (anonymous):

the answer is 9.15 am

OpenStudy (anonymous):

how you find 142 ??

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

I added.... I dont get this.....